### v1

# Nonflat bands and chiral symmetry in magic-angle twisted bilayer graphene

Abstract
In this work, we derive effective theories for magic-angle twisted bilayer graphene (MATBG) based on the mean-field normal state at charge neutrality of a tight-binding model, thereby including the renormalizations coming from integrating out high-energy modes. In these models, the flat bands display a sizable increase of the bandwidth, and the corresponding wavefunctions flow towards the limit of perfect particle-hole symmetry and sublattice polarization. We further represent the flat bands in the ’vortex Chern’ basis and also discuss the implications on the dynamics regarding the ’flat’ and ’chiral’ symmetries of MATBG, as manifested in the symmetry-broken states at neutrality.
preprint: APS/123-QED
Introduction.
Magic-angle twisted bilayer graphene (MATBG), as a type of moiré material [1, 2], hosts a mesoscopic unit cell and flat electronic bands at charge neutrality. It represents a tunable platform for the study of strong interactions and topology [3, 4, 5], as well as their interplay, thus attracting great attention in the recent years. Experimental and theoretical efforts have unveiled a rich phase diagram, exhibiting in a single material phenomena like correlated states [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], Kondo physics [28, 29, 30, 31, 32, 33, 34] and superconductivity [35, 36, 17, 18, 19, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], among others [47, 48, 49, 50, 51].
The large number of atoms in the moiré unit cell, and consequently large number of electronic bands, calls for the construction of effective theories of MATBG with a small number of degrees of freedom, that ought to capture accurately the dispersion, symmetry and topology of the low-energy bands. Among such theories there are continuum models [52, 53, 54, 55], including the seminal Bistritzer-MacDonald (BM) model [56], Wannier orbital [57, 58, 59, 60, 61, 62, 63, 64, 65], momentum space [66, 67] and the topological heavy fermion (THF) model [68].
Their very small bandwidth renders the flat bands prone to distortions coming from otherwise minor perturbations, like lattice relaxation [54, 53, 69, 70] or strain [69, 71, 72]. In this respect, an often overlooked piece of the physics is the interacting effect of the high-energy degrees of freedom removed in the effective theory, but nonetheless implicit in the ground state. We argue that virtual processes between the ’remote’ (high-energy) and ’active’ (low-energy) modes are significant and must be accounted for [73], at least approximately. An earlier work addressed this problem using a Renormalization Group approach [74].
In this Letter, we construct a normal state at charge neutrality that preserves the symmetries and incorporates the Hartree-Fock potentials of the occupied graphene $\pi$ orbitals self-consistently; any low-energy theory of MATBG ought to reproduce the dispersion and wavefunctions of the normal, rather than the non-interacting, state. In this work we obtained effective models for MATBG using a many-body projection scheme onto the central bands of the normal state, sketched in Fig. 1.
In the renormalized system the dispersion widens drastically, with the flat bandwidth reaching several times the non-interacting one and becoming comparable to the energy scale of the Coulomb interactions among flat electrons. Furthermore, we find that the flat-band Hilbert space flows towards the limit of perfect particle-hole symmetry and sublattice polarization.
At the same time, there are important modifications of the symmetry of MATBG [6, 75, 68]. The $U(4)_{\text{flat}}$ symmetry of the ’nonchiral-flat’ limit, that relies on the existence of very flat bands, is strongly broken; instead, the ’chiral-nonflat’ $U(4)_{\text{chiral}}$ group prevails as an approximate symmetry independently of the bandwidth. Such properties become apparent in the symmetry-broken phases. Lastly, we comment on the experimental imprints of our results on the recently observed spiral states [76].
Figure 1: Many-body projection scheme. First, we obtain a renormalized, self-consistent, symmetry-preserving state. We then project onto excitations of the central bands, resulting in a renormalized dispersion, $\mathcal{H}_{0}$, with nonflat bands.
Figure 2: (a) Band structure of the symmetric state for different interaction strengths, compared to the non-interacting (bare) bands. (b) Hilbert space of the bare and renormalized ($\epsilon_{r}=10,$ $U=4$ eV) flat bands. In the top row we show the positive eigenvalue of the sublattice operator, $\overline{\sigma}_{z}$, (sublattice polarization). The remaining eigenvalues are obtained from $180^{\circ}$ rotations and sign changes. In the bottom rows we plot the two distinct singular values of the projected particle-hole operator $C_{2z}\overline{P}$, which are degenerate in pairs.
Symmetric state and many-body projection.
We start with a tight-binding model for MATBG with one orbital per carbon atom. The twist angle is chosen to $\theta=1.05^{\circ}$, resulting in a unit cell with 11908 atoms, and in-plane atomic relaxations are included following Ref. [77]. The crystalline symmetries of the lattice include six-fold rotations around the $z$ axis, $C_{6z}$, and two-fold rotations around the $y$ axis between the layers, $C_{2y}$. Below we will be addressing the products $C_{3z}=C_{6z}^{2}$, $C_{2z}=C_{6z}^{3}$ and $C_{2x}=C_{2y}C_{2z}$ also. Spinless time-reversal, $\mathcal{T}$, acting as complex conjugation, is a symmetry if we neglect spin-orbit coupling.
[…]
The total Hamiltonian, ${H}$, consists of the tight-binding part and the interaction, ${H}=H_{\text{TB}}+{H}_{\text{int}}$, with
$H_{\text{TB}}=$ $\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s},$
$H_{\text{int}}=$ $\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\delta n_{\bm{r}s}\delta n_{\bm{r^{\prime}}s^{\prime}}+\sum_{\bm{r}}U\delta n_{\bm{r}\uparrow}\delta n_{\bm{r}\downarrow}.$ (1)
Here, $c^{\dagger}_{\bm{r}s}$($c_{\bm{r}s}$) is the creation (annihilation) operator of an electron with spin $s$ at position $\bm{r}$ and $t(\bm{r})$ denotes the hopping function [80, 81]; $\delta n_{\bm{r}s}=n_{\bm{r}s}-\frac{1}{2}=c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\frac{1}{2}$ is the spin-$s$ density at position $\bm{r}$ relative to a uniform background.
We find a self-consistent solution of ${H}$ in mean-field (Hartree-Fock) theory that preserves spin rotations, time-reversal symmetry and the crystallographic symmetries (see the Supplementary Material 111see Supplementary Material, which includes the references [94, 77, 95, 80, 81, 90, 96, 75, 6, 85, 4, 5, 84, 54, 69, 87, 88, 68, 97, 92, 15, 41, 98, 99, 100, 101, 102, 103] for details). In Fig. 2(a) we plot the band structures of the symmetric state with $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0.5$ eV (the weak coupling limit corresponds to the physical situation if the self-screening is strong [83]), and compare them with the bare band structure. The flat bandwidth increases approximately linearly with $\epsilon_{r}^{-1}$, growing from $8.7$ meV to $15.3$ meV to $48.4$ meV in the bare, $\epsilon_{r}=50$ and $\epsilon_{r}=10$ bands, respectively. Another relevant quantity is the gap between the flat and remote bands, controlling the validity of the flat-band projection [66]. The gaps to the remote bands increase from $20.2$ to $26.8$ meV on the electron side and from $23.9$ to $30.8$ meV on the hole side in the bare and $\epsilon_{r}=10$ bands, respectively.
Not only the energies, but also the wave functions are renormalized. We study the properties of Hilbert space of the flat bands in Fig. 2(b). We plot the sublattice polarization, this is, the eigenvalues of the sublattice operator projected to the flat bands, $\overline{\sigma}_{z}$, and the singular values of the ’particle-hole’ operator [84, 5, 54, 69] projected to the flat bands, $C_{2z}\overline{P}$. The system flows towards a perfectly sublattice polarized and particle-hole symmetric limit with eigenvalues (singular values) equal to $1$.
Using the projector onto many-body excitations of the lowest $n_{B}$ bands of the normal state, $\mathcal{P}$, we obtain an effective model for MATBG with Hamitonian ${\mathcal{H}}=\mathcal{P}H\mathcal{P}=\mathcal{H}_{0}+{\mathcal{H}}_{\text{int}}$,
[…]
${\mathcal{H}}_{\text{int}}=\frac{1}{2}{\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}}V(\bm{r}-\bm{r^{\prime}})\delta{\rho}_{\bm{r}s}\delta{\rho}_{\bm{r^{\prime}}s^{\prime}}+\sum_{\bm{r}}U\delta{\rho}_{\bm{r}\uparrow}\delta{\rho}_{\bm{r}\downarrow}.$ (2)
In this expression, the operator $c^{\dagger}_{\bm{k}ns}(c_{\bm{k}ns})$ creates (annihilates) an electron with momentum $\bm{k}$, band $n$ and spin $s$ in the mean-field basis, and $n,m$ are restricted to the central $n_{B}$ bands, $n,m=\{-n_{B}/2,...,n_{B}/2\}$. In both cases, the $n_{B}$ lowest bands are gapped from the remaining bands at all $\bm{k}$, so the projector is well-defined.
The matrix ${h}(\bm{k})_{nm}$ contains the dispersion of the normal state plus a counter-term to avoid double-counting the valence bands $20<n\leq-n_{B}/$ 2, and a term related to the order of the fermionic operators, see the Supplementary Materials [82] for details. On the other hand, ${\rho}_{\bm{r}s}$ is the density relative to charge neutrality,
$\delta{\rho}_{\bm{r}s}={\rho}_{\bm{r}s}-{\rho}_{0\bm{r}},\quad\quad{\rho}_{\bm{r}s}=\sum_{\mathclap{\bm{k}\bm{q}nm}}\langle\bm{k}n|\bm{r}\rangle\langle\bm{r}|\bm{q}m\rangle c^{\dagger}_{\bm{k}ns}c_{\bm{q}ms},$
${\rho}_{0\bm{r}}=\sum_{\mathclap{\bm{k},n<0}}|\langle\bm{r}|\bm{k}n\rangle|^{2}\approx\frac{1}{2}\sum_{\bm{k}n}|\langle\bm{r}|\bm{k}n\rangle|^{2},$ (3)
where $\langle\bm{r}|\bm{k}n\rangle$ is the wave function amplitude of mode $c^{\dagger}_{\bm{k}ns}$ at position $\bm{r}$ and again $n,m=\{-n_{B}/2,...,n_{B}/2\}$. The choice of $\rho_{0\bm{r}}$ is consistent with the approximate particle-hole symmetry.
The band structure of $\mathcal{H}_{0}$ is shown in Fig. 3(a) with $\epsilon_{r}=10$, $U=4$ eV, for $n_{B}=4$ and $20$. In both cases, the dispersion is only slightly narrower than in the normal state.
Vortex Chern gauge and nonflat bands.
In the flat-band ($n_{B}=4$) theory, we obtain a valley and sublattice polarized basis by diagonalizing the projeccted valley ($\overline{\tau}_{z}$) [85] and sublattice ($\overline{\sigma}_{z}$) operators. We use $\tau_{i}$ and $\sigma_{i}$ to denote the identity ($i=0$) and Pauli ($i=x,y,z$), matrices in valley and sublattice space, valley $K(K^{\prime})$ and sublattice $A(B)$ having eigenvalue 1(1)111(-1)1 ( - 1 ) under $\tau_{z}$ and $\sigma_{z}$, respectively. The resulting bands are topological with Chern numbers equal to $\sigma_{z}\tau_{z}$ [6, 4, 5].
The ’Chern gauge’ [6, 75] fixes at each $\bm{k}$ the relative phase between bands with the same Chern number, leaving an ambiguity of the form $\text{exp}(i\phi(\bm{k})\sigma_{z}\tau_{z})$ [6, 86, 13] (see the Supplementary Materials [82] for details on the gauge fixing and symmetry representations). We fix $\phi(\bm{k})$ by requiring that the Wannier functions obtained from the Chern bands are maximally localized [87, 88], following the algorithm of Ref. [87]; the Bloch states are periodic and smooth except at the $\Gamma_{M}$ point, where the Berry connection acquires a vortex with a winding number equal to the Chern number. This ’vortex Chern gauge’ is defined up to a global sign of all wave functions.
The renormalized dispersion can now be decomposed into Hermitian components,
[…]
where $c^{\dagger}_{\bm{k}}=c^{\dagger}_{\bm{k}\alpha}$ is a vector of creation operators and ${h}(\bm{k})={h}(\bm{k})_{\alpha\beta}$ a matrix, $\alpha,\beta$ denoting the valley-sublattice-spin flavors and $i,j=0,x,y,z$. By construction ${h}(\bm{k})$ is proportional to the unit spin matrix, and by $C_{2z}\mathcal{T}$ and valley conservation symmetries only the coefficients ${h}_{00}(\bm{k})$, ${h}_{0z}(\bm{k})$, ${h}_{x0}(\bm{k})$, ${h}_{yz}(\bm{k})$, ${h}_{y0}(\bm{k})$ and ${h}_{xz}(\bm{k})$ are nonzero.
In Fig. 3(b) we show the non-vanishing coefficients for $\epsilon_{r}=10$, $U=4$ eV. The dominant terms are ${h}_{x0}$ and ${h}_{yz}$, ${h}_{00}$ is also sizable at $\Gamma_{M}$. The ’minimal’ Hamiltonian with ${h}(\bm{k})={h}_{x0}(\bm{k})\sigma_{x}+{h}_{yz}(\bm{k})\sigma_{y}\tau_{z}$ inherits the Dirac cones at $K,K^{\prime}$ and has a bandwidth of $41.9$ meV, to be compared with the bare bandwidth of $8.7$ meV. Furthermore, the representation $\text{exp}(-2\pi i/3\ \sigma_{z}\tau_{z})$ of $C_{3z}$ forces the phase of ${h}_{x0}(\bm{k})+i{h}_{yz}(\bm{k})$ to wind $2$ mod $3$ times around $\Gamma_{M}$. Band topology manifests itself at $\Gamma_{M}$ where the representation of $C_{3z}$ changes to $\mathbb{1}$, hence ${h}_{x0}(\Gamma_{M})+i{h}_{yz}(\Gamma_{M})$ need not vanish and there is a vortex in ${h}_{x0}(\bm{k})+i{h}_{yz}(\bm{k})$.
In Fig. 3(c) we summarize the main features of ${h(\bm{k})}$ as a function of $1/\epsilon_{r}$ The different components grow linearly with $1/\epsilon_{r}$, ${h}_{x0}$ and ${h}_{yz}$ becoming the largest with increasing coupling, already prevailing at $\epsilon_{r}\lesssim 50$.
Figure 3: (a) Comparison of the bare bands, the bands of the symmetric state and the bands of the projected models with $n_{B}=4$ and $20$. (b) The nonzero components of ${h}(\bm{k})$ in the vortex Chern basis for $n_{B}=4$. Notice the different scale bars ($\epsilon_{r}=10$, $U=4$ eV in (a-b)). (c) The components of ${h}(\bm{k})$ as a function of $1/\epsilon_{r}$; data points for $\epsilon_{r}^{-1}=0$, $U=0$ eV, $\epsilon_{r}=50$, $U=0.5$ eV and $\epsilon_{r}=10$, $U=4$ eV. The average over the Brillouin zone is denoted avBZ and the lines are linear fits.
Symmetry and correlated states.
Here, we comment on the consequences of our renormalization procedure on the symmetry in the $n_{B}=4$ system. For a detailed analysis of the symmetries of MATBG, including larger $n_{B}$, we refer the reader to the Supplementary Materials [82] as well as to Refs. [75, 6, 68].
The most important outcome is the widening of the flat bands by the large component ${h}_{x0}(\bm{k})\sigma_{x}+{h}_{yz}(\bm{k})\sigma_{y}\tau_{z}$. The generator $\mathcal{G}_{\text{flat}}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}\sigma_{y}\tau_{x}c_{\bm{k}}$ is then broken and the system diverges from the $U(4)_{\text{flat}}$ symmetry of the ’nonchiral-flat’ limit. On the other hand, $\mathcal{G}_{\text{chiral}}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}\sigma_{x}\tau_{y}c_{\bm{k}}$ commutes with this dominant term thus the ’chiral-nonflat’ $U(4)_{\text{chiral}}$ symmetry is approximately preserved.
With respect to the interaction, the breaking of $\mathcal{G}_{\text{chiral}}$ and $\mathcal{G}_{\text{flat}}$ can be assessed by the departure from one of the eigenvalues and singular values of $\overline{\sigma}_{z}$ and $C_{2z}\overline{P}$, respectively, displayed in Fig. 2(b). They increase in the symmetric state, thus enhancing the symmetry of ${\mathcal{H}}_{\text{int}}$; this phenomenon was already noted in Ref. [74] for $\overline{\sigma}_{z}$.
In order to investigate the interacting phases, we have performed Hartree-Fock calculations in the flat-band theory. The self-consistent states can be described by the one-particle density matrix, $P(\bm{k})_{\alpha\beta}=\langle c^{\dagger}_{\bm{k}\alpha}c_{\bm{k}\beta}\rangle=\frac{1}{2}\Big{(}\mathbb{1}_{\alpha\beta}+Q(\bm{k})_{\alpha\beta}\Big{)}$.
We study several symmetry-breaking phases at charge neutrality: the Kramers intervalley coherent state (KIVC, $Q(\bm{k})=\sigma_{y}\tau_{y}$), orbital polarized (OP, $Q(\bm{k})=\sigma_{z}$, also called valley Hall), valley polarized (VP, $Q(\bm{k})=\tau_{z}$), time-reversal invariant intervalley coherent (TIVC, $Q(\bm{k})=\sigma_{x}\tau_{x}$) [7, 6, 13], Quantum Anomalous Hall (QAH, $Q(\bm{k})=\sigma_{z}\tau_{z}$) [12, 13, 89, 6, 7], nematic semimetal (NSM, $Q(\bm{k})=\cos(\pi/3)\sigma_{x}-\sin(\pi/3)\sigma_{y}\tau_{z}$) [12, 11, 10, 89] and the spin polarized state (SP, $Q(\bm{k})=s_{z}$, the spin-$z$ matrix). The states above serve as initial seeds in the self-consistency loop; for VP we enforced $C_{2z}\mathcal{T}$ and valley conservation symmetries, and for TIVC and NSM, $C_{2z}$ and $\mathcal{T}$. The energies of the converged states are reported in Table 1.
Invariably in all phases, the kinetic energy penalty induces $Q(\bm{k})\sim\sigma_{x},\sigma_{y}\tau_{z}$ around the $\Gamma_{M}$ point, coming from the lower bands of $\mathcal{H}_{0}$. This is apparent in Fig. 4(b) where we plot the $Q(\bm{k})$ matrix of KIVC and OP. Concurrently, the wide dispersion enforces TIVC, VP and SP to be gapless, while KIVC, OP and QAH are insulating, see Fig. 4(a). These (semi)metallic states are destabilized by $\mathcal{H}_{0}$ as compared to the others. Moreover, the approximate chiral symmetry, $|\text{KIVC}\rangle\approx\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{OP}\rangle$, $|\text{VP}\rangle\approx\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{TIVC}\rangle$ is manifest in Fig. 4(a) and (b).
$E-E_{\text{SYM}}$ (meV)
$\epsilon_{r}=10$, $U=4$ eV $\epsilon_{r}=50$, $U=0.5$ eV
KIVC $-14.171$ $-3.579$
QAH $-12.808$ $-3.295$
OP $-11.994$ $-3.252$
SP $-11.483$ $-2.823$
NSM $-9.835$ $-2.812$
VPL $-9.075$ $-2.526$
TIVC $-7.872$ $-2.407$
Table 1: Condensation energy of the self-consistent states at charge neutrality relative to the energy of the symmetry-preserving normal state.
Figure 4: (a) Band structure of the Kramers intervalley coherent (KIVC), orbital polarized (OP), time-reversal invariant intervalley coherent (TIVC) and valley polarized (VP) states. The approximate chiral symmetry relating KIVC and OP, and TIVC and VP is manifest. (b) Order parameters, defined by $Q(\bm{k})=\sum_{ij}\langle\sigma_{i}\tau_{j}(\bm{k})\rangle\sigma_{i}\tau_{j}$, of the KIVC and OP states. The quantities $\langle\sigma_{y}\tau_{y}(\bm{k})\rangle$ and $\langle\sigma_{z}(\bm{k})\rangle$ are related by a chiral rotation, and the quantities $\langle\sigma_{x}(\bm{k})\rangle$ and $\langle\sigma_{y}\tau_{z}(\bm{k})\rangle$ are trivial under chiral rotations corresponding to filling the valence bands of $\mathcal{H}_{0}$. The coupling strength is $\epsilon_{r}=10$, $U=4$ eV.
Discussion.
In this work, we performed large scale (11908 atoms per unit cell) computations in an atomistic model of MATBG and obtained a mean-field symmetric state at charge neutrality. In this state, the dispersion and the wave functions are renormalized such that the electrons are decoupled at the one-loop (Hartree-Fock) level. We argue that effective models with a reduced bandwidth reproducing this symmetric phase take account of the many-body effects coming from the removed degrees of freedom, as opposed to models obtained from a naive truncation of the bare Hilbert space.
We have constructed effective theories using a many-body projection onto the lowest $4$ and $20$ bands per spin of the symmetric state. The $4$-band model corresponds to the flat band projection, and the $20$-band model describes an energy window of about ±$\pm 250$ meV. In these theories the flat bands widen drastically, with a bandwidth comparable to the interaction scale, demonstrating the relevance of the renormalizations induced by integrating out the remote bands
The widening of the flat bands is directly related to the breaking of the $U(4)_{\text{flat}}$ symmetry. Instead, we find that the chiral group $U(4)_{\text{chiral}}$ is an approximate symmetry of the system, broken at a scale of some tens of meV$/\epsilon_{r}$.
In a previous study [74], the authors studied the BM model with an energy cutoff of order of the graphene bandwidth, and integrated out the high-energy bands via a Renormalization Group procedure. However, the BM Hamiltonian is only accurate within a certain energy range, making the procedure uncontrolled above a threshold. In contrast, our scheme is self-consistent rather than perturbative, and more importantly, our tight-binding model accurately describes the wavefunctions within the full bandwidth of the $\pi$ orbitals. This explains the discrepancies in the final outcomes, for instance, the band widening in [74] is relatively weak.
On another note, our ordering of the Coulomb interaction (Eq. 1) is consistent with the ’average’ subtraction scheme in MATBG literature [75, 74]. Another widely used protocol is the ’graphene’ subtraction scheme [6, 9], where $\mathcal{H}_{\text{int}}$ is normal-ordered [90] with respect to the ground state of two decoupled layers at neutrality.
It has been assumed that due to the subtraction, the normal state is close to the bare ground state in the graphene scheme [6, 12, 11, 13, 89, 91]. Contrary to this expectation, our calculations with graphene subtraction, found in the Supplementary Materials [82], show significant band widening also; the flat bands reach a bandwidth of $43.8$ meV for $\epsilon_{r}=10$. More notably, we have identified a topological phase transition on the flat bands at some value of $\epsilon_{r}$ between $50$ and $10$.
The results presented here promise to have extensive consequences in our understanding of MATBG. We already mentioned the pattern of explicit symmetry breaking, $U(4)\times U(4)\to U(4)_{\text{chiral}}$ instead of $U(4)\times U(4)\to U(4)_{\text{flat}}$ which is conventionally assumed [68, 33, 32, 15]. In close connection to this, the VP and TIVC states are ’dynamically enforced’ to be gapless in contrast to most theory results [68, 92, 15].
In another regard, let us consider the strain-stabilized, moiré translations-breaking, inconmensurate Kekulé spiral (IKS) phase [86, 89, 14, 76]. The IKS momentum, $\bm{q}_{\text{IKS}}$, is determined by a nesting condition in the band structure of the nematic semimetal (NSM), and it can be measured in STM experiments [93]. The NSM can be significantly modified by the band renormalization, for instance the high energy cost if the Dirac nodes lie close to $\Gamma_{M}$ from $\mathcal{H}_{0}$ competes with the exchange gain if the Dirac cones lie at the BZ center [10]. Thus, a comparison between the theoretically predicted and the experimentally determined $\bm{q}_{\text{IKS}}$ provides a direct test of our findings.
Acknowledgements.
This work was supported by grant PID2020–113164 GBI00 funded by MCIN/AEI/10.13039/501100011033, PID2023-146461NB-I00 and PRE2021-097070 funded by Ministerio de Ciencia, Innovación y Universidades, and the CSIC Research Platform on Quantum Technologies PTI-001. The access to computational resources of CESGA (Centro de Supercomputación de Galicia) is also gratefully acknowledged.
[…]
Supplementary Material for ”Nonflat bands and chiral symmetry in magic-angle twisted bilayer graphene”
I Lattice geometry and tight-binding model
In a graphene monolayer the lattice vectors are $\bm{a}_{1}=a_{0}(1/2,\sqrt{3}/2)$ and $\bm{a}_{2}=a_{0}(-1/2,\sqrt{3}/2)$, with $a_{0}=2.46$ the lattice constant. The atoms on the latice positions set up the $A$ sublattice, and the atoms displaced from the $A$ atoms by $(\bm{a}_{1}+\bm{a_{2}})/3$ set up the $B$ sublattice. The dual lattice vectors are then $\bm{G}_{1}=4\pi/\sqrt{3}a_{0}(\sqrt{3}/2,1/2)$ and $\bm{G}_{2}=4\pi/\sqrt{3}a_{0}(-\sqrt{3}/2,1/2)$. The band structure displays the characteristic Dirac cones around the two $K$ points, $K=(\bm{G_{2}}-\bm{G_{1}})/3$ and $K^{\prime}=-K$. Our convention differs from other works in which $K=(\bm{G_{1}}-\bm{G_{2}})/3$ by a $180^{\circ}$ rotation, or equivalently by an interchange of the $A$ and $B$ sublattice labels.
We set up the twisted bilayer geometry starting from a graphene bilayer with the two layers stacked on top of each other, at vertical positions $z=\pm d_{0}/2=\pm 1.6755$. The top layer gets rotated by an angle $\theta/2$ and the bottom layer by $-\theta/2$, with the center of rotation being the center of one of the carbon hexagons. For a generic twist angle $\theta$, the resulting structure is not strictly periodic, only for particular angles there is a commensurate crystalline structure. A subset of such commensurate angles can be parameterized as $\cos(\theta)=1-1/(6n^{2}+6n+2)$ with integer $n$ [94]. The unit vectors of the resulting superlattice are $\bm{L}_{1}=L_{M}(\sqrt{3}/2,1/2)$ and $\bm{L}_{2}=L_{M}(-\sqrt{3}/2,1/2)$ with $L_{M}=a_{0}\sqrt{3n^{2}+3n+1}$ the moiré lattice constant. The reciprocal lattice in turn has unit vectors $\bm{g}_{1}=2\pi/\sqrt{3}L_{M}(1/2,\sqrt{3}/2)$ and $\bm{g}_{2}=2\pi/\sqrt{3}L_{M}(1/2,\sqrt{3}/2)$.
The magic-angle of twisted bilayer graphene sits between $1.0^{\circ}$ and $1.1^{\circ}$ in different experiments and theoretical models. In this work we chose $\theta=1.05^{\circ}$ ($n=31$) such that the bandwidth of the flat bands of our tight-binding model is minimal. The moiré lattice constant is of $13.4$ nm and the number of atoms in the unit cell is $11908$. We have included the effects of in-plane lattice relaxation, i.e. symmetry-preserving intrinsic strain, using the model of Ref. [77].
In the bilayer, the crystallographic symmetry of $6$-fold rotations around the twist axis (the $z$ axis through one of the hexagons), denoted $C_{6z}$, is preserved. Yet the mirror symmetry about the $y$ axis of each individual layer is lost in the twisted structure, as it relates a lattice with twist angle $\theta$ to one with angle $-\theta$. Instead, there is a related symmetry consisting of $180^{\circ}$ rotations around the $y$ axis at $z=0$, denoted $C_{2y}$. This operation acts as a mirror symmetry about the $y$ axis plus an interchange of the layers. The group elements $C_{6z}^{2}$, $C_{6z}^{3}$ and $C_{2z}C_{2y}$ performing $3$-fold $z$-rotations, $2$-fold $z$-rotations and $180^{\circ}$ rotations around the $x$ axis are denoted $C_{3z}$, $C_{2z}$ and $C_{2x}$, respectively. Furthermore, the spin-orbit coupling in carbon is negligible, so there is an additional spinless time-reversal symmetry, $\mathcal{T}$, acting as complex conjugation.
We set up a microscopic tight-binding model for the carbon $\pi$ orbitals of MATBG. We employ the Slater-Koster [95] hopping integral given in Refs. [80, 81]. The hopping function depends on the distance between atoms and the angle of the radius vector with the $z$ axis,
$\begin{gathered}t(\bm{r})=-V_{pp\pi}(r)\big{(}1-\cos^{2}(\varphi_{\bm{r}})\big{)}+V_{pp\sigma}(r)\cos^{2}(\varphi_{\bm{r}}),\\ V_{pp\pi}(r)=V_{pp\pi}^{0}\ \text{exp}\big{(}-(r-a_{cc})/r_{0}\big{)},\quad\quad\quad V_{pp\sigma}(r)=V_{pp\sigma}^{0}\ \text{exp}\big{(}-(r-d_{0})/r_{0}\big{)},\end{gathered}$ (S1)
where $\bm{r}=(x,y,z)$, $r=\sqrt{x^{2}+y^{2}+z^{2}}$ and $\cos(\varphi_{\bm{r}})=z/r$. The constants are $a_{cc}=a_{0}/\sqrt{3}$ (the carbon-carbon distance), $V^{0}_{ppp\pi}=2.7$ eV, $V^{0}_{pp\sigma}=0.48$ eV and $r_{0}=0.0453$ nm.
Figure S1: (a) Top view of MATBG with a twist angle of $6.01^{\circ}$ ($n=5$). The lattice vectors $\bm{L}_{1}$, $\bm{L}_{2}$ and the moiré Wigner-Seitz cell are marked. (b) Triangular loops used for computing the valley charge. (c) Brillouin zones of the rotated top (blue) and bottom (orange) monolayers, and the hexagonal Brillouin zones of the coupled layers (black), drawn in an extended scheme. The $K(K^{\prime})$ points of the original graphenes are depicted. (d) Moiré Brillouin zone of MATBG, with the high-symmetry $\Gamma_{M}$, $M_{M}$, $K_{M}$, $K^{\prime}_{M}$ points and the reciprocal lattice vectors $\bm{g}_{1}$, $\bm{g}_{2}$ labeled.
II Mean-field decoupling and many-body projection
Let us recall the full interacting tight-binding Hamiltonian,
$\begin{split}&{H}=H_{\text{TB}}+{H}_{\text{int}}\\ &H_{\text{TB}}=\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}\\ &{H}_{\text{int}}=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\bigg{(}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\frac{1}{2}\bigg{)}\bigg{(}c^{\dagger}_{\bm{r^{\prime}}s^{\prime}}c_{\bm{r^{\prime}}s^{\prime}}-\frac{1}{2}\bigg{)}+\frac{1}{2}\sum_{\bm{r}s}U\bigg{(}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\frac{1}{2}\bigg{)}\bigg{(}c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}-\frac{1}{2}\bigg{)},\end{split}$ (S2)
with $V(\bm{r}-\bm{r^{\prime}})=\frac{14.4\text{eV}\times\text{\AA}}{\epsilon_{r}}\sum_{n=-\infty}^{\infty}\frac{(-1)^{n}}{||\bm{r}-\bm{r^{\prime}}+100\text{\AA}\times n\bm{\hat{z}}||}$ and $\bar{s}$ the spin projection opposite to $s$. We solve $H$ using the standard mean-field decoupling [90],
${H}\approx{H}_{\text{MF}}=$ $\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\bigg{(}\langle c^{\dagger}_{\bm{r}s^{\prime}}c_{\bm{r}s^{\prime}}\rangle_{0}-\frac{1}{2}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}$
$-\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r})\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{0}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\sum_{\bm{r}s}U\bigg{(}\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{0}-\frac{1}{2}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant},$ (S3)
were $\langle...\rangle_{0}$ denotes the expectation value in a state of reference. We look for a self-consistent solution where the reference state corresponds to the ground state of ${H}_{\text{MF}}$. We assume a spin-symmetric state, hence $\langle c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}\bar{s}}\rangle_{0}=0$ and the spins are decoupled in the Fock channel (third term).
In practice, we assume periodic boundary conditions and preserved moiré translations, and work in the basis of Bloch waves,
$c^{\dagger}_{\bm{k}\bm{\delta}s}=\frac{1}{\sqrt{N_{c}}}\sum_{\bm{R}_{\ell}}e^{i\bm{k}\cdot(\bm{\delta}+\bm{R}_{\ell})}c^{\dagger}_{\bm{\delta}+\bm{R}_{\ell},s},$ (S4)
with $N_{c}$ the number of unit cells and $\bm{k}$ is the Bloch momentum inside the rhombic Brillouin zone, $\bm{k}=(n_{1}/\sqrt{N_{c}})\bm{G_{1}}+(n_{2}/\sqrt{N_{c}})\bm{G_{2}}$ for integers $n_{1},n_{2}=0,...,\sqrt{N_{c}}-1$. The index $\bm{\delta}$ denotes an atomic position inside the Wigner-Seitz cell, and $\bm{R}_{\ell}$ denotes the lattice vectors. Different Bloch momenta are coupled in ${H}$, but by translation symmetry they become decoupled in the mean-field Hamiltonian ${H}_{\text{MF}}$. The mean-field basis $c^{\dagger}_{\bm{k}ns}$ diagonalizes ${H}_{\text{MF}}$,
$\begin{split}{H}_{\text{MF}}=&\sum_{\bm{k}ns}\varepsilon^{\text{MF}}_{n}(\bm{k})c^{\dagger}_{\bm{k}ns}c_{\bm{k}ns}+\text{constant},\\ c^{\dagger}_{\bm{k}ns}=&\sum_{\bm{\delta}}\langle\bm{k}\bm{\delta}|\bm{k}n\rangle c^{\dagger}_{\bm{k}\bm{\delta}s}=\sum_{\bm{r}}\langle\bm{r}|\bm{k}n\rangle c^{\dagger}_{\bm{r}s},\end{split}$ (S5)
where $\langle\bm{k}\bm{\delta}|\bm{k}n\rangle$ and $\langle\bm{r}|\bm{k}n\rangle$ are the wave function amplitudes (independent of spin) of mode $c^{\dagger}_{\bm{k}ns}$ in the Bloch and real-space basis, respectively. The conduction bands have band index $n=1,2,...$ with increasing $n$ corresponding to increasing energy, and $n=-1,-2,...$ denotes the valence bands with increasing $|n|$ corresponding to decreasing energy. In the self-consistent algorithm we used the ground state of the bare system ($H_{\text{TB}}$) as initial seed and imposed the crystallographic symmetries and spinless time-reversal symmetry at each iteration. We used a $6\times 6$ discretization of the Brillouin zone and the ODA algorithm [96] to accelerate convergence.
With the renormalized symmetric state at hand, we wish to construct an effective model that accounts for the relevant low-energy excitations accurately. In order to achieve this, we assume that the mean-field bands with $n<-n_{B}/2$ are fully occupied and with $n>n_{B}/2$ fully empty, and we allow for fluctuations in the central $n_{B}$ bands. We restrict to such states via the the many-body projector
$\begin{split}\mathcal{P}=&\sum_{I}|\Psi_{I}\rangle\big{\langle}\Psi_{I}|,\\ |\Psi_{I}\rangle=&\prod_{\bm{k}ns\in I}c^{\dagger}_{\bm{k}ns}\prod_{\bm{k}s,n<{-n_{B}}/{2}}c^{\dagger}_{\bm{k}ns}|0\rangle,\end{split}$ (S6)
where $I$ denotes the set of indices of one particular element of a complete basis of the subspace of many-body excitations of the central $n_{B}$ bands, and the sum over $I$ runs over all $2^{N_{C}\times n_{B}\times 2}$ basis elements. In other words, $\mathcal{P}$ projects to all $n$-body ($n=1,2,...,N_{c}\times n_{B}\times 2$) excitations on the central $n_{B}$ bands.
Consider a generic two-body Hamiltonian $V=\frac{1}{2}\sum_{abcd}V_{abcd}c^{\dagger}_{a}c^{\dagger}_{b}c_{c}c_{d}$. (in our case, $V=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c^{\dagger}_{\bm{r^{\prime}}s^{\prime}}c_{\bm{r^{\prime}}s^{\prime}}c_{\bm{r}s}+\frac{1}{2}\sum_{\bm{r}s}Uc^{\dagger}_{\bm{r}s}c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}c_{\bm{r}s}$). We can deduce the form of $\mathcal{P}V\mathcal{P}$ by considering the action of $V$ on a generic two-particle state in the central bands. We will work in the mean-field basis and use $a,b,c,d$ for generic modes, $\alpha,\beta,\gamma,\delta$ for modes within the central bands and $i,j$ for modes with $n<-n_{B}/2$. After some algebra, we get
$\frac{1}{2}\sum_{abcd}V_{abcd}c^{\dagger}_{a}c^{\dagger}_{b}c_{c}c_{d}\ c_{\alpha}^{\dagger}c_{\beta}^{\dagger}\prod_{j}c^{\dagger}_{j}\big{|}0\big{\rangle}=$ $\frac{1}{2}\sum_{\gamma\delta}(V_{\gamma\delta\beta\alpha}-V_{\gamma\delta\alpha\beta})c_{\gamma}^{\dagger}c_{\delta}^{\dagger}\prod_{j}c^{\dagger}_{j}\big{|}0\big{\rangle}$
$+\sum_{\gamma i}(V_{\gamma ii\alpha}-V_{i\gamma i\alpha})c_{\gamma}^{\dagger}c_{\beta}^{\dagger}\prod_{j}c^{\dagger}_{j}\big{|}0\big{\rangle}-\sum_{\gamma i}(V_{\gamma ii\beta}-V_{i\gamma i\beta})c_{\gamma}^{\dagger}c_{\alpha}^{\dagger}\prod_{j}c^{\dagger}_{j}\big{|}0\big{\rangle}$
$+\frac{1}{2}\sum_{ik}(V_{ikki}-V_{ikik})c^{\dagger}_{\alpha}c^{\dagger}_{\beta}\prod_{j}c^{\dagger}_{j}\big{|}0\big{\rangle}+\text{others},$ (S7)
where again we have restricted the final states to be of the form $c_{\alpha}^{\dagger}c_{\beta}^{\dagger}\prod_{j}c^{\dagger}_{j}|0\rangle$; different states are collectively denoted as ’others’. Hence, the many-body projected $V$ reads
$\mathcal{P}V\mathcal{P}=\frac{1}{2}\sum_{\alpha\beta\gamma\delta}V_{\alpha\beta\gamma\delta}c^{\dagger}_{\alpha}c^{\dagger}_{\beta}c_{\gamma}c_{\delta}+\sum_{\alpha\beta}\sum_{i}(V_{\alpha ii\beta}-V_{i\alpha i\beta})c^{\dagger}_{\alpha}c_{\beta}+\frac{1}{2}\sum_{ik}(V_{ikki}-V_{ikik})$ (S8)
Using Eq. (S7) one can check that the expression in Eq. (S8) has the same action as $V$ if we restrict to initial and final states contained in $\mathcal{P}$. One now sees that, crucially, the occupied modes that have been integrated out induce new terms which are precisely the Hartree and Fock potentials acting on the central bands. The constant piece is the energy of the projected-out Fermi sea.
On the other hand, on the operator $O=\sum_{ab}O_{ab}c^{\dagger}_{a}c_{b}$ the projection truncates to the $\alpha$ modes and adds a constant,
$\mathcal{P}O\mathcal{P}=\sum_{\alpha\beta}O_{\alpha\beta}c^{\dagger}_{\alpha}c_{\beta}+\sum_{i}O_{ii}.$ (S9)
We have just deduced that the effective model with Hamiltonian $\mathcal{P}{H}\mathcal{P}$ accounts for the many-body effects of the integrated valence bands, $n<-n_{B}/2$, via the one-particle terms in $\mathcal{P}{H}_{\text{int}}\mathcal{P}$. Now, remember that the dispersion of the symmetric state $\varepsilon^{\text{MF}}_{n}(\bm{k})$ already includes the mean-field potentials from the valence bands $n<0$, hence the projected Hamiltonian can be written in terms of the dispersion $\varepsilon^{\text{MF}}_{n}(\bm{k})$ plus a counter term from the valence bands with $20>n\geq-n_{B}/$ 2 to avoid double counting,
${\mathcal{H}}=\mathcal{P}{H}\mathcal{P}=$ $\sum_{\bm{k}ns}\varepsilon^{\text{MF}}_{n}(\bm{k})c^{\dagger}_{\bm{k}ns}c_{\bm{k}ns}+\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\ d^{\dagger}_{\bm{r}s}d^{\dagger}_{\bm{r^{\prime}}s^{\prime}}d_{\bm{r^{\prime}}s^{\prime}}d_{\bm{r}s}+\sum_{\bm{r}}U\ d^{\dagger}_{\bm{r}\uparrow}d^{\dagger}_{\bm{r}\downarrow}d_{\bm{r}\downarrow}d_{\bm{r}\uparrow}$
$+\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r^{\prime}})\bigg{(}\sideset{}{{}^{\prime}}{\sum}_{\bm{q},m<0}\langle\bm{r}|\bm{q}m\rangle\langle\bm{q}m|\bm{r^{\prime}}\rangle\bigg{)}d^{\dagger}_{\bm{r}s}d_{\bm{r^{\prime}}s}$
$-2\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r^{\prime}})\rho_{0\bm{r^{\prime}}}\ d^{\dagger}_{\bm{r}s}d_{\bm{r}s}-\sum_{\bm{r}}U\rho_{0\bm{r}}\ d^{\dagger}_{\bm{r}s}d_{\bm{r}s}$
[…]
with
$d^{\dagger}_{\bm{r}s}={\sum_{\bm{k}n}}^{\prime}\langle\bm{k}n|\bm{r}\rangle c^{\dagger}_{\bm{k}ns},\quad\quad\quad\rho_{0\bm{r}}=\sideset{}{{}^{\prime}}{\sum}_{\bm{k},n<0}|\langle\bm{r}|\bm{k}n\rangle|^{2}.$ (S11)
The primed summations indicate that $n$ is restricted to the central bands, $n=-n_{B}/2,...,n_{B}/2$. The Fock and Hartree counter-terms are written in the second and third line, respectively. The ’constant’ is the zero-point energy of the frozen Fermi sea.
Now, the truncated position operators satisfy the relations $\{d^{\dagger}_{\bm{r}s},d^{\dagger}_{\bm{r^{\prime}}s^{\prime}}\}=\{d_{\bm{r}s},d_{\bm{r^{\prime}}s^{\prime}}\}=0$, $\{d^{\dagger}_{\bm{r}s},d_{\bm{r^{\prime}}s^{\prime}}\}=\sum_{\bm{k}n}^{\prime}\langle\bm{r^{\prime}}|\bm{k}n\rangle\langle\bm{k}n|\bm{r}\rangle\delta_{ss^{\prime}}$. In order to write the interacting term in density-density form ($\sim d^{\dagger}_{\bm{r}s}d_{\bm{r}s}d^{\dagger}_{\bm{r^{\prime}}s^{\prime}}d_{\bm{r^{\prime}}s^{\prime}}$) we must take account of the anomalous anticommutator when changing the order of operators. In sum, the projected Hamiltonian reads
$\begin{split}&{\mathcal{H}}={\mathcal{H}}_{0}+{\mathcal{H}}_{\text{int}}+\text{constant},\\ &{\mathcal{H}}_{0}={\sum_{\bm{k}ns}}^{\prime}{h}(\bm{k})_{nm}c^{\dagger}_{\bm{k}ns}c_{\bm{k}ms},\\ &{\mathcal{H}}_{\text{int}}=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\delta\rho_{\bm{r}s}\delta\rho_{\bm{r^{\prime}}s^{\prime}}+\sum_{\bm{r}}U\delta\rho_{\bm{r}\uparrow}\delta\rho_{\bm{r}\downarrow}.\end{split}$ (S12)
The exchange components have been combined with $\varepsilon^{\text{MF}}_{n}(\bm{k})$ in ${\mathcal{H}}_{0}$, while the Hartree components are implicit in ${\mathcal{H}}_{\text{int}}$ through the background density $\rho_{0\bm{r}}$,
$\begin{gathered}h(\bm{k})_{nm}=\varepsilon^{\text{MF}}_{n}(\bm{k})\delta_{nm}+\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}}}V(\bm{r}-\bm{r^{\prime}})\langle\bm{k}n|\bm{r}\rangle\langle\bm{r^{\prime}}|\bm{k}m\rangle\bigg{(}\sideset{\,}{{}^{\prime}}{\sum}_{\bm{q},\ell<0}\langle\bm{r}|\bm{q}\ell\rangle\langle\bm{q}\ell|\bm{r^{\prime}}\rangle-\sideset{}{{}^{\prime}}{\sum}_{\bm{q},\ell>0}\langle\bm{r}|\bm{q}\ell\rangle\langle\bm{q}\ell|\bm{r^{\prime}}\rangle\bigg{)},\\ \delta\rho_{\bm{r}s}=\rho_{\bm{r}s}-\rho_{0\bm{r}},\quad\quad\quad\quad\quad\rho_{\bm{r}s}=d^{\dagger}_{\bm{r}s}d_{\bm{r}s}.\end{gathered}$ (S13)
This is the form of the effective model studied in the main text.
III Vortex Chern basis of the flat bands
In this section we describe a gauge-fixing procedure on the flat-band manifold. We follow the procedure of Ref. [6], constructing the ’sublattice basis’, and use the conventions of Ref. [75] for the definition and gauge fixing of the ’particle-hole’ operator, see below. This ’Chern gauge’ was originally defined for the bare flat bands, but we can equally impose it for the central bands of the mean-field state because both systems are adiabatically connected. We assume a spinless system; the physical system with spin includes two identical copies of the basis below.
Consider the microscopic valley ($\tilde{\tau}_{z}$) [85] and sublattice ($\tilde{\sigma}_{z}$) operators,
$\begin{split}\tilde{\tau}_{z}&=\frac{i}{3\sqrt{3}}\Bigg{(}\sum_{\begin{subarray}{c}\ll\bm{r},\bm{r^{\prime}}\gg,s\\ \bm{r},\bm{r^{\prime}}\in\text{sl}\ A\end{subarray}}\eta_{\bm{rr^{\prime}}}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}-\sum_{\begin{subarray}{c}\ll\bm{r},\bm{r^{\prime}}\gg,s\\ \bm{r},\bm{r^{\prime}}\in\text{sl}\ B\end{subarray}}\eta_{\bm{rr^{\prime}}}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}\Bigg{)},\\ \tilde{\sigma}_{z}&=\sum_{\bm{r}\in\text{sl}\ A,s}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\sum_{\bm{r}\in\text{sl}\ B,s}c^{\dagger}_{\bm{r}s}c_{\bm{r}s},\end{split}$ (S14)
where $\ll\bm{r},\bm{r^{\prime}}\gg$ denotes next-nearest neighbors, ’sl’ is short for sublattice and $\eta_{\bm{rr^{\prime}}}=\pm 1$ depending if the pair $\bm{r},\bm{r^{\prime}}$ is part of a clockwise or counterclockwise triangle, see Fig. S1(b). Both operators conserve the Bloch momentum and $[\tilde{\tau}_{z},\tilde{\sigma}_{z}]=0$. The valley operator is constructed so that graphene plane waves near the $K$($K^{\prime}$) point have expectation values close to $1$($-1$).
Here and throughout, single-particle operators restricted to the flat bands will be indicated by a line, $\overline{\phantom{o}}$’ (Eq. (S9) would then read $\mathcal{P}O\mathcal{P}=\overline{O}+\sum_{i}O_{ii}$). We start by simultaneously diagonalizing the flat-band matrices at each $\bm{k}$, $[\overline{\sigma}_{z}(\bm{k})]_{nm}=\langle\bm{k}n|\tilde{\sigma}_{z}|\bm{k}m\rangle$ and $[\overline{\tau}_{z}(\bm{k})]_{nm}=\langle\bm{k}n|\tilde{\tau}_{z}|\bm{k}m\rangle$. Valley $K$($K^{\prime}$) states are identified as the states with positive (negative) eigenvalue under $\overline{\tau}_{z}$. We use $\tau_{i}$, $\sigma_{i}$ for the identity ($i=0$) and Pauli ($i=x,y,z$) matrices acting in the valley and sublattice polarized sectors, respectively. Valley $K(K^{\prime})$ has eigenvalue 1(1)111(-1)1 ( - 1 ) under $\tau_{z}$, and we assign eigenvalue 1(1)111(-1)1 ( - 1 ) to sublattice $A(B)$ under $\sigma_{z}$. The resulting bands are topological with Chern numbers equal to $\sigma_{z}\tau_{z}$ [6, 4, 5].
The ’particle-hole’ operator, $\tilde{P}$, is emergent in the continuum limit of MATBG [84, 5, 54, 69]. In practice, we consider the product $C_{2z}\tilde{P}$ rather than $\tilde{P}$ (understand the product of single-particle operators in second quantization as the second-quantized version of the product of the corresponding first-quantized operators). For ease of notation we will not distinguish between the microscopic, indicated with $\tilde{\phantom{o}}$’, and projected, indicated with $\overline{\phantom{o}}$’, versions of the crystallographic and time-reversal symmetries; their distinction must be understood within the context. Note moreover that the exact symmetries are block-diagonal in the band basis. For the definition of $C_{2z}\tilde{P}$, see Section IV.
The symmetry $C_{2z}\mathcal{T}$ and the particle-hole operator are local in Bloch momentum. We impose the ’Chern gauge’ on the valley-sublattice basis, defined by the following form of $C_{2z}\mathcal{T}$ and $C_{2z}\overline{P}$,
$\begin{split}C_{2z}\mathcal{T}=\sigma_{x}\mathcal{K},\\ C_{2z}\overline{P}=\sigma_{y}\tau_{x},\end{split}$ (S15)
with $\mathcal{K}$ denoting the antilinear complex conjugation operator. This gauge respects the relations $\{C_{2z}\overline{P},\tau_{z}\}=[C_{2z}\overline{P},C_{2z}\mathcal{T}]=\{C_{2z}\overline{P},\overline{\sigma}_{z}\}=[C_{2z}\mathcal{T},\tau_{z}]=\{C_{2z}\mathcal{T},\overline{\sigma}_{z}\}=0$.
As a matter of fact, the flat Hilbert space is not invariant under the action of the particle-hole operator (in other words, $C_{2z}\tilde{P}$ does not commute with the projector to the flat bands), forbidding the representation $C_{2z}\overline{P}=\sigma_{y}\tau_{x}$. This is evidenced by the singular vales of $C_{2z}\overline{P}$ in Fig. S2(a) and Table S2 (see also Refs. [54, 104]). Instead, we replace $C_{2z}\overline{P}$ by the unitary matrix $C_{2z}P=C_{2z}\overline{P}/\big{(}C_{2z}\overline{P}(C_{2z}\overline{P})^{\dagger}\big{)}^{1/2}$ and enforce Eq. (S15) for $C_{2z}P$. A similar procedure was applied when we assigned eigenvalues ±$\pm 1$ under $\sigma_{z}$ to the sublattice polarized states even though the eigenvalues of $\overline{\sigma}_{z}$ are different from ±$\pm 1$ (Fig. S2(a) and Table S2). Notice that because $C_{2z}P\neq C_{2z}\overline{P}$, $\{C_{2z}\overline{P},\overline{\sigma}_{z}\}$ need not be equal to 00 and $C_{2z}P$ might display components of the form $\tau_{x}$, $\tau_{y}$ and $\sigma_{z}\tau_{y}$ that cannot be removed with a gauge choice. Nevertheless, we find that these terms are negligible and we can satisfy Eq. (S15) up to numerical precision.
Two-fold rotations and time-reversal symmetries invert the momentum, $\bm{k}\to-\bm{k}$, and three-fold rotations rotate the momentum by $120^{\circ}$, $\bm{k}\to C_{3z}(\bm{k})$, up to translations by a dual lattice vector. We impose their action on the flavor degrees of freedom to be
$\begin{split}C_{2z}&=\begin{cases}\sigma_{x}\tau_{x}\quad&\bm{k}=\Gamma_{M}\\ -\sigma_{x}\tau_{x}\quad&\bm{k}\neq\Gamma_{M}\end{cases},\\ \mathcal{T}&=\begin{cases}\tau_{x}\mathcal{K}\quad&\bm{k}=\Gamma_{M}\\ -\tau_{x}\mathcal{K}\quad&\bm{k}\neq\Gamma_{M}\end{cases},\\ C_{3z}&=\begin{cases}\mathbb{1}\quad&\bm{k}=\Gamma_{M}\\ \text{exp}(-2\pi i/3\ \sigma_{z}\tau_{z})\quad&\bm{k}\neq\Gamma_{M}\end{cases}.\end{split}$ (S16)
The departure at $\Gamma_{M}$ from a ’uniform’ representation can be traced back to the topological nature of the Chern bands.
As is, the Chern gauge defined by Eqs. (S15),(S16) exhibits an ambiguity consisting of rotations of the states with the same Chern number, $\text{exp}(i\phi(\bm{k})\sigma_{z}\tau_{z})$. We fix this phase by requiring that the Wannier functions obtained from the Chern bands are maximally localized [87, 88], following the algorithm of Ref. [87]. By $C_{2z}\mathcal{T}$, the maximal localization procedure yields the same phase $\phi_{K}(\bm{k})$ for the $KA$ and $KB$ polarized bands, and $\phi_{K^{\prime}}(\bm{k})$ for the $K^{\prime}A$ and $K^{\prime}B$ bands. Given that there exists some particle-hole asymmetry, i.e. $C_{2z}\overline{P}\neq C_{2z}P$, $\phi_{K}(\bm{k})$ need not be equal to $\phi_{K^{\prime}}(\bm{k})$. In practice we find that they differ by at most $1^{\circ}$. We allow for this small error and take $\phi(\bm{k})$ to be the average. With this gauge transformation, the Bloch states are periodic and smooth except at the $\Gamma_{M}$ point, where the Berry connection acquires a vortex with a winding number equal to the Chern number [87, 88].
Finally, we fix a global phase by imposing $C_{2y}\mathcal{T}=\mathcal{K}$ at $\Gamma_{M}$; our ’vortex Chern gauge’ is then defined up to a global sign of all wave functions.
As a final comment, we have found that the phase $\phi(\bm{k})$ satisfies $\phi(\bm{k})=\phi(-\bm{k})$ exactly, and $\phi(\bm{k})\approx\phi(C_{3z}(\bm{k}))$ only approximately (we attribute this to having taken the average of $\phi_{K}(\bm{k})$ and $\phi_{K^{\prime}}(\bm{k})$ before). Consequently, the action of $C_{3z}$ is slightly modified to $C_{3z}=\text{exp}(-\lambda(\bm{k})\sigma_{z}\tau_{z})$ with $\lambda(\Gamma_{M})=0$, $\lambda(K_{M})=\lambda(K_{M}^{\prime})=-2\pi/3$ and $\lambda(\bm{k})\approx-2\pi/3$ elsewhere, with deviations form $-2\pi/3$ always smaller than $3^{\circ}$.
IV Symmetry of MATBG
In the following we study the internal (flavor) symmetries of MATBG. For alternative discussions as well as additional details, see Refs. [6, 75, 68].
In the spirit of the continuum models, we can decompose a generic wave function into valley components,
$\langle\bm{r}|\psi\rangle=\sum_{\tau}e^{i\tau\bm{K}_{l}\cdot\bm{r}}f^{\psi}_{\tau\sigma l}(\bm{r}).$ (S17)
The phases $e^{\pm i\bm{K}_{l}\cdot\bm{r}}$ oscillate rapidly on the graphene lattice, whereas the envelope functions $f_{\tau\sigma l}(\bm{r})$ vary slowly on the graphene lattice and depend on the valleys $\tau=\pm$ and on the sublattice $\sigma=A,B$ and layer $l=\text{top}\ (t),\text{bottom}\ (b)$ of the atom at $\bm{r}$. $(-)\bm{K}_{l}$ is the $K\ (K^{\prime})$ point of layer $l$, see Fig. S1. We assume that $|\psi\rangle$ has a definite spin projection that is left implicit. This decomposition is only justified for states at low energies coming from the valleys of the parent graphene monolayers. In other words, the decomposition can be written only in effective models for MATBG where remote bands above a certain cutoff have been integrated out. In continuous models, the functions $f$ defined at the atomic positions are promoted to smooth functions in the continuum.
Symmetry of ${\mathcal{H}}_{\text{int}}$.
Making use of Eq. (S17), consider the matrix element of $\mathcal{H}_{\text{int}}$ for generic states $\alpha,\beta,\gamma,\delta$ with valleys $\eta_{\alpha},\eta_{\beta},\eta_{\gamma},\eta_{\delta}$,
${H}_{\text{int},\alpha\beta\gamma\delta}=$ $\sum_{\bm{r}\neq\bm{r^{\prime}}}V(\bm{r}-\bm{r^{\prime}})e^{i(\tau_{\delta}-\tau_{\alpha})\bm{K}_{l}\cdot\bm{r}}e^{i(\tau_{\gamma}-\tau_{\beta})\bm{K}_{l^{\prime}}\cdot\bm{r^{\prime}}}f^{\alpha}_{\tau_{\alpha}\sigma l}(\bm{r})^{*}f^{\delta}_{\tau_{\delta}\sigma l}(\bm{r})f^{\beta}_{\tau_{\beta}\sigma^{\prime}l^{\prime}}(\bm{r^{\prime}})^{*}f^{\gamma}_{\tau_{\gamma}\sigma^{\prime}l^{\prime}}(\bm{r^{\prime}})$
$+U\sum_{\bm{r}}e^{i(\tau_{\delta}+\tau_{\gamma}-\tau_{\alpha}-\tau_{\beta})\bm{K}_{l}\cdot\bm{r}}f^{\alpha}_{\tau_{\alpha}\sigma l}(\bm{r})^{*}f^{\delta}_{\tau_{\delta}\sigma l}(\bm{r})f^{\beta}_{\tau_{\beta}\sigma l}(\bm{r})^{*}f^{\gamma}_{\tau_{\gamma}\sigma l}(\bm{r}).$ (S18)
The nature of the Coulomb interaction forces the implicit spin components $s_{\alpha}=s_{\delta}$, $s_{\beta}=s_{\gamma}$ for both the long range and Hubbard terms, and $s_{\alpha}=\bar{s}_{\beta}$ for the Hubbard term. Notice also that we have omitted an essentially trivial piece coming from the $-\frac{1}{2}$ subtraction, see Eq. (S2).
Clearly, the matrix element of Eq. (S18) is the same as ${\mathcal{H}}_{\text{int},\alpha\beta\gamma\delta}$ if the initial and final states belong to the central bands. Nonetheless, ${\mathcal{H}}_{\text{int}}$ includes also a nontrivial local potential
$\mathcal{H}_{\text{int}}\supset\sum_{\bm{r}s}\pi(\bm{r})\rho_{\bm{r}s},\quad\quad\quad\quad\quad\pi(\bm{r})=-2\sum_{\bm{r^{\prime}}\neq\bm{r}}V(\bm{r}-\bm{r^{\prime}})\rho_{0\bm{r^{\prime}}}-U\rho_{0\bm{r}},$ (S19)
see Eqs. (S12), (S13), that we ought to consider separately.
Valley-spin flavor symmetries.
Looking at the first line of Eq. (S18), only if $\tau_{\alpha}=\tau_{\delta}$ and $\tau_{\beta}=\tau_{\gamma}$ the matrix elements is non vanishing. Indeed, if $\tau_{\delta}=-\tau_{\alpha}$, we can take the $\bm{r}$, $\bm{r}+\bm{a}_{1}$ and $\bm{r}+\bm{a}_{2}$ and compute the phases $e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}$, $e^{2i\tau_{\delta}\bm{K}_{l}\cdot(\bm{r}+\bm{a}_{1}/2)}=e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}e^{\pm 2\pi i/3}$ and $e^{2i\tau_{\delta}\bm{K}_{l}\cdot(\bm{r}+\bm{a}_{2}/2)}=e^{2i\tau_{\delta}\bm{K}_{l}\cdot\bm{r}}e^{\mp 2\pi i/3}$. Given that the $f$ functions and the potential $V(\bm{r})$ vary slowly on atomic distances, $V(\bm{r}+\bm{a}_{1,2})\approx V(\bm{r})$, $f(\bm{r}+\bm{a}_{1,2})\approx f(\bm{r})$, the sum over the corresponding three atoms in ${\mathcal{H}}_{\text{int},\alpha\beta\gamma\delta}$ (modulo lattice relaxation and twist) cancels by $01+e^{\pm 2\pi i/3}+e^{\mp 2\pi i/3}=$ 0. Similarly occurs if $\tau_{\gamma}=-\tau_{\beta}$.
The matrix element then has the structure $\mathcal{H}_{\text{int},\alpha\beta\gamma\delta}\sim\delta_{\tau_{\alpha}\tau_{\delta}}\delta_{\tau_{\beta}\tau_{\gamma}}\delta_{s_{\alpha}s_{\delta}}\delta_{s_{\beta}s_{\gamma}}$ with some coefficient that does not depend on the spin. Thus, the long range part of ${\mathcal{H}}_{\text{int}}$ enjoys a $SU(2)_{K}\times SU(2)_{K^{\prime}}\times U(1)_{\text{V}}$ symmetry. The two $SU(2)$ correspond to independent spin rotations in each valley, and $U(1)_{\text{V}}$ to the valley charge conservation enforcing $\tau_{\alpha}+\tau_{\beta}=\tau_{\gamma}+\tau_{\delta}$. A heuristic explanation for the $SU(2)_{K}\times SU(2)_{K^{\prime}}$ symmetry is that different valleys only interact via the total density in the Hartree channel, and because valley-dependent spin rotations do not change the density from each valley, these generate a symmetry.
On the other hand, the potential $\pi(\bm{r})$, being smooth on atomic distances, cannot produce high-momentum transfer scatterings between the two valleys, and it is independent of spin. Consequently, the symmetry is exactly preserved.
Combined with the $U(1)_{\text{C}}$ of electric charge conservation we have a $U(2)_{K}\times U(2)_{K^{\prime}}$ group with the 8 generators
$\big{[}s_{i},{\mathcal{H}}_{\text{int}}\big{]}=\big{[}s_{i}\tau_{z},{\mathcal{H}}_{\text{int}}\big{]}=0,\quad\quad\quad\quad\quad(i=0,x,y,z).$ (S20)
The operators $s_{i}$ are the identity ($i=0$) and Pauli ($i=x,y,z$) spin operators.
With respect to the second line of Eq. (S18), by a similar argument we deduce that valley charge is conserved, $\tau_{\alpha}+\tau_{\beta}=\tau_{\gamma}+\tau_{\delta}$. However, only global spin rotations preserve the Hubbard Hamiltonian, breaking $U(2)_{K}\times U(2)_{K^{\prime}}$ down to the physical symmetry of $SU(2)_{\text{spin}}\times U(1)_{\text{V}}$. Furthermore, the long range interaction also breaks $U(2)_{K}\times U(2)_{K^{\prime}}$. Indeed, in the argument presented previously we assumed that the the potential $V(\bm{r})$ is slowly varying on atomic distances. This is not true near $\bm{r}=0$ where there is a steep $1/|\bm{r}|$ dependence. In turn, there are symmetry breaking components coming from small distances. $U(2)_{K}\times U(2)_{K^{\prime}}$-breaking interactions are called ’intervalley Hund’s’ couplings.
The electron-phonon interaction generates antiferromagnetic Hund’s couplings (i.e. stabilizing opposite spins on opposite valleys) that can be decisive in selecting the ground state [97, 92, 15, 41, 98, 99]. On the other hand, the long range interaction is Hund-antiferromagentic [100] and the on-site Hubbard interaction, ferromagnetic [41, 101]. These energy scales are much smaller than the chiral and flat-breaking terms discussed below; we leave their study for a future publication.
Chiral symmetry.
Now, consider the ’chiral’ operator $\tilde{\sigma}_{z}$ in Eq. (S14). The exponential $\text{exp}(i\varphi\tilde{\sigma}_{z})$ attaches the phase $e^{i\varphi}$ to $f_{\tau Al}$ and $e^{-i\varphi}$ to $f_{\tau Bl}$. Clearly, this is an exact symmetry of ${{H}}_{\text{int}}$ because each phase encounters the complex conjugate in the matrix element, Eq. (S18).
Nonetheless, the mode given by $\text{exp}(i\varphi\tilde{\sigma}_{z})|\alpha\rangle$ generically does not belong to the subspace of central bands (in other words, it is not possible in general to construct a perfectly sublattice polarized basis in the effective theory), spoiling the chiral symmetry in $\mathcal{H}_{\text{int}}$
To overcome this problem, we redefine the sublattice operator on the central bands as $\sigma_{z}=\sum_{\alpha\beta}[\sigma_{z}]_{\alpha\beta}c^{\dagger}_{\alpha}c_{\beta}$, with the matrix $\sigma_{z}=\overline{\sigma}_{z}/\big{(}\overline{\sigma}_{z}\overline{\sigma}_{z}^{\dagger}\big{)}^{1/2}$ and $[\overline{\sigma}_{z}]_{\alpha\beta}=\langle\alpha|\tilde{\sigma}_{z}|\beta\rangle$. This $\sigma_{z}$ here is the same as $\sigma_{z}$ acting on the Chern basis defined below Eq. (S14), having the property that it is the unitary matrix closest to $\overline{\sigma}_{z}$ (in the norm $||M||=\text{tr}(M^{\dagger}M)$). We can assess the discrepancy between $\overline{\sigma}_{z}$ and $\sigma_{z}$ via the singular values or, equivalently up to a sign, the eigenvalues of $\overline{\sigma}_{z}$, see Table (S2). Singular values equal to $1$ imply that $\sigma_{z}=\overline{\sigma}_{z}$.
With respect to the local potential, in the perfectly sublattice polarized limit the symmetry is exact owing to $\sigma_{z}\rho_{\bm{r}z}\sigma_{z}=\rho_{\bm{r}s}$. However, from the fact that $\overline{\sigma}_{z}\neq\sigma_{z}$ there appear intersublattice scatterings introducing some amount of symmetry breaking.
In turn, the exact chiral symmetry of ${H}_{\text{int}}$ turns into an approximate symmetry in the low-energy theory,
$\big{[}\sigma_{z}\tau_{z},{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (S21)
the degree of non-symmetry indicated by the sublattice polarization of the central bands (we have attached the factor $\tau_{z}$ of the exact valley charge conservation, without need to change the argument).
Particle-hole symmetry.
Let us turn to the particle-hole operator, $C_{2z}\tilde{P}$. This is a single-particle operator, local in real space and defined by its action on the $f$ functions in Eq. (S17),
$C_{2z}\tilde{P}=\sigma_{x}\tau_{y}\mu_{y},$ (S22)
with $\mu_{y}$ the Pauli $y$ matrix in layer space, the top (bottom) layer having eigenvalue 1(1)111\ (-1)1 ( - 1 ) under $\mu_{z}$. This operator is only properly defined on the continuous wave functions, but not in the atomistic theory. In practice, we obtain the smooth $f$ functions by interpolation in a valley polarized basis, and later they are sampled in the atomic positions of the opposite sublattice and layer. The inner products of the transformed wave functions with the basis wave functions define the matrix elements of $C_{2z}\overline{P}$.
We will argue now that the matrix element ${{H}}_{\text{int},\alpha^{\prime}\beta^{\prime}\gamma^{\prime}\delta^{\prime}}$ between modes $\eta^{\prime}=C_{2z}\tilde{P}|\eta\rangle$ ($\eta=\alpha,\beta,\gamma,\delta$) is approximately equal to ${{H}}_{\text{int},\alpha\beta\gamma\delta}$. For that, we may establish a one-to-one correspondence between atoms $\bm{r}$ and $\bm{s}$, with $\bm{s}$ belonging to the opposite sublattice and layer to $\bm{r}$ and with approximately the same in-plane coordinates. Then, we can replace the sums over $\bm{r},\bm{r^{\prime}}$ in Eq. (S18) by sums over $\bm{s},\bm{s^{\prime}}$ and conclude that, in the $U(2)_{K}\times U(2)_{K^{\prime}}$ limit where $\tau_{\alpha}=\tau_{\delta}$, $\tau_{\beta}=\tau_{\gamma}$ and the valley phases cancel, ${{H}}_{\text{int},\alpha\beta\gamma\delta}={{H}}_{\text{int},\alpha^{\prime}\beta^{\prime}\gamma^{\prime}\delta^{\prime}}$ to a very good approximation. This property hints to the existence of a continuous symmetry generated by $C_{2z}\tilde{P}$. Indeed, in Appendix D of Ref. [102] we showed that $[{H}_{\text{int}},C_{2z}\tilde{P}]=0$ (assuming $U(2)_{K}\times U(2)_{K^{\prime}}$ and up to effects from the discrete lattice), also we provide more details on the implementation of $C_{2z}\tilde{P}$ in the atomistic model in Appendix C of Ref. [102].
Again, the symmetry is not well defined in the effective theory hence, analogously to $\sigma_{z}$, we define $C_{2z}P=\sum_{\alpha\beta}[C_{2z}P]_{\alpha\beta}c^{\dagger}_{\alpha}c_{\beta}$, with the matrix $C_{2z}P=C_{2z}\overline{P}/\big{(}C_{2z}\overline{P}(C_{2z}\overline{P})^{\dagger}\big{)}^{1/2}$ and $[C_{2z}\overline{P}]_{\alpha\beta}=\langle\alpha|C_{2z}\tilde{P}|\beta\rangle$.
With regards to the local potential, in the case of perfect particle-hole symmetry, and neglecting intervalley scatterings, we have $C_{2z}{P}\rho_{\bm{r}s}C_{2z}{P}=\rho_{\bm{s}s}$. Then, provided $\rho_{0\bm{r}}=\rho_{0\bm{s}}$, the potential respects the $C_{2z}P$ symmetry. We find that $\rho_{0\bm{r}}\approx\rho_{0\bm{s}}$ to a good approximation, as shown in Fig. S2(b); moreover we also find that $\rho_{0\bm{r}}\approx\frac{1}{2}\sum_{\bm{k}n}^{\prime}|\langle\bm{r}|\bm{k}n\rangle|^{2}$. These properties of $\rho_{0\bm{r}}$ can be partly understood from the constrains imposed by the crystallographic and valley symmetries and the properties of the normal state under $C_{2z}P$ and $\sigma_{z}$, see Eq. (IV) (where we can replace $\mathcal{H}_{0}$ by the normal state, $H_{\text{MF}}$). Notice also that in $\pi(\bm{r})$ the density is convoluted with the Coulomb potential, which tends to smoothen the inhomogeneities making the relation $\pi(\bm{r})\approx\pi(\bm{s})$ tighter. Particle-hole asymmetry, $C_{2z}\overline{P}\neq C_{2z}P$, translates to some amount of symmetry breaking.
We conclude that $C_{2z}P$ generates an approximate symmetry in the low-energy theory,
$\big{[}C_{2z}P,{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (S23)
and the strength of symmetry breaking can be assessed by the singular values of $C_{2z}\overline{P}$, displayed in Fig. S2(a) and Table (S2).
$R$ symmetry.
Consider as well the product of operators
$\tilde{R}=C_{2z}\tilde{P}\tilde{\sigma}_{z}\tau_{z}=\sigma_{y}\tau_{x}\mu_{y},$ (S24)
which is clearly conserved by the interaction, $[{H}_{\text{int}},\tilde{R}]=0$, being the product of exact symmetries. Again, the symmetry becomes approximate in the effective theory with generator $R$,
$\big{[}R,{\mathcal{H}}_{\text{int}}\big{]}\approx 0,$ (S25)
and the strength of the symmetry breaking can be deduced from the singular values of $\overline{R}$. Here we have defined $R$ and $\overline{R}$ analogously to $\overline{\sigma}_{z}$ and $C_{2z}\overline{P}$, and $\sigma_{z}$ and $C_{2z}P$, respectively.
It is important to note that $\overline{R}$ need not be equal to $C_{2z}\overline{P}\overline{\sigma}_{z}\tau_{z}$; the projection of $\tilde{R}=C_{2z}\tilde{P}\tilde{\sigma}_{z}\tau_{z}$ is not the same as the product of the projections of $C_{2z}\tilde{P}$ and $\tilde{\sigma}_{z}\tau_{z}$. Nonetheless, the low-energy Hilbert space is very particle-hole symmetric (the singular values of $C_{2z}\overline{P}$ in the flat bands are of the order of $0.99$), thus we conclude that in fact $\overline{R}\approx C_{2z}\overline{P}\overline{\sigma}_{z}\tau_{z}$. Consequently, the singular values of $\overline{R}$ will be approximately equal to the singular values of $\overline{\sigma}_{z}$, of the order of $0.5-0.9$ in the flat bands.
By the same reason, the unitary matrix $R$ is not necessarily equal to $C_{2z}{P}\sigma_{z}\tau_{z}$. For instance, in the flat-band theory with the Chern gauge, the relations $\{R,\tau_{z}\}=\{R,C_{2z}\mathcal{T}\}=\{R,\tilde{\sigma}_{z}\}=0$ impose the form $R=\sin(\nu(\bm{k}))\sigma_{y}\tau_{y}+\cos(\nu(\bm{k}))\sigma_{x}\tau_{y}$ for some momentum-dependent angle $\nu(\bm{k})$. Observe that the property $[C_{2z}\tilde{P},\tilde{R}]=0$ no longer holds for $[C_{2z}P,R]$. Again, almost particle-hole symmetry obliges $R\approx$ $C_{2z}{P}\sigma_{z}\tau_{z}=\sigma_{x}\tau_{y}$, i.e. $\nu(\bm{k})\approx 0$. Indeed, $\nu(\bm{k})$ is at most ±$\pm 1.2^{\circ}$. We ignore such small deviations from $\nu(\bm{k})=0$ and consider the symmetry as generated by $R=C_{2z}{P}\sigma_{z}\tau_{z}=\sigma_{x}\tau_{y}$.
[…]
Full symmetry group: chiral and flat limits
Once identified the ’hidden’ symmetries, we can obtain the full symmetry group by commuting the $U(2)_{K}\times U(2)_{K^{\prime}}$ generators of Eq. (S20) with $C_{2z}P$, $R$ and/or $\sigma_{z}\tau_{z}$. For simplicity, let us consider the spinless system, in which $U(2)_{K}\times U(2)_{K^{\prime}}$ turns to $U(1)_{\text{C}}\times U(1)_{\text{V}}$, generated by the identity, $\mathbb{1}$, and $\tau_{z}$. Let us also work in the Chern basis of the $n_{B}=4$ model for concreteness. The symmetry generators are written as $\mathcal{G}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}Gc_{\bm{k}}$ with $c^{\dagger}_{\bm{k}}=\big{(}c^{\dagger}_{\bm{k}\alpha}\big{)}$ a vector and $G=[G]_{\alpha\beta}$ a matrix in the Chern basis; let us work with just the $G$ matrices for neatness.
With $C_{2z}P$, we can obtain $S_{x}=\frac{-i}{2}[\tau_{z},C_{2z}P]=\sigma_{x}\tau_{x}$. Denoting $S_{0}=\mathbb{1}$, $S_{y}=C_{2z}P$ and $S_{z}=\tau_{z}$ we see that $\{S_{0},S_{x},S_{y},S_{z}\}$ generates a $U(2)$ group. On the other hand, with the $R$ symmetry we get a different $U(2)$ group with $S^{\prime}_{0}=\mathbb{1}$, $S^{\prime}_{x}=R$, $S^{\prime}_{y}=\frac{i}{2}[\tau_{z},R]=\sigma_{y}\tau_{y}$ and $S^{\prime}_{z}=\tau_{z}$ (the $\sigma_{z}\tau_{z}$ symmetry results unphysical, spoiled by both $\mathcal{H}_{\text{int}}$ and $\mathcal{H}_{0}$, and we avoid discussing it). In the limit with both $C_{2z}P$ and $R$ (hence also $\sigma_{z}\tau_{z}$), there is a total of $8$ generators, $\{\mathbb{1},\sigma_{x}\tau_{x},\sigma_{x}\tau_{y},\sigma_{y}\tau_{x}\sigma_{y}\tau_{y},\tau_{z},\sigma_{z}\tau_{z}\}$. One can check that an alternative basis of the Lie algebra is $P_{\pm}\{\mathbb{1},\sigma_{x}\tau_{x},\sigma_{x}\tau_{y},\tau_{z}\}P_{\pm}$, with $P_{\pm}=1/2(\mathbb{1}\pm\sigma_{z}\tau_{z})$ the projector to the Chern $\pm$ sector. It becomes clear that this is the Lie algebra of a $U(2)\times U(2)$ group of independent $U(2)$ transformations in each Chern sector.
The same procedure can be followed straightforwardly including the spin; the $U(2)$ groups are enlarged to $U(4)$ and the $U(2)\times U(2)$ to $U(4)\times U(4)$. The different limits and their symmetry groups are listed in Table (S1). The case with $U(2)_{K}\times U(2)_{K^{\prime}}$ and $C_{2z}P$ is dubbed ’flat’ limit and the symmetry $U(4)_{\text{flat}}$ (the dispersion $\mathcal{H}_{0}$ is prone to breaking this symmetry, as we discuss below). On the other hand, with $U(2)_{K}\times U(2)_{K^{\prime}}$ and $R$ we are in the ’chiral-nonflat’ limit and the symmetry is denoted $U(4)_{\text{chiral}}$. Note that in the main text we use the notations $C_{2z}P=\mathcal{G}_{\text{flat}}$ and $R=\mathcal{G}_{\text{chiral}}$.
symmetry group generators $\mathcal{G}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}Gc_{\bm{k}}$
chiral-flat $U(4)\times U(4)$ $G=\{t_{a},t_{a}\sigma_{x}\tau_{y},t_{a}\sigma_{y}\tau_{x},t_{a}\sigma_{z}\}$
chiral-nonflat $U(4)_{\text{chiral}}$ $G=\{t_{a},t_{a}\sigma_{y}\tau_{x}\}$
nonchiral-flat $U(4)_{\text{flat}}$ $G=\{t_{a},t_{a}\sigma_{x}\tau_{y}\}$
nonchiral-nonflat $U(2)_{K}\times U(2)_{K^{\prime}}$ $G=\{t_{a}\}=\{s_{i},s_{i}\tau_{z}\}\quad(i=0,x,y,z)$
physical $SU(2)_{\text{spin}}\times U(1)_{\text{V}}\times U(1)_{\text{C}}$ $G=\{s_{i},\tau_{z},\mathbb{1}\}\quad(i=x,y,z)$
Table S1: Different limits and their symmetry groups in the Chern basis of $n_{B}=4$ MATBG. The explicit symmetry breaking hierarchy is $U(4)\times U(4)\to U(4)_{\text{chiral}}\to U(2)_{K}\times U(2)_{K^{\prime}}\to SU(2)_{\text{spin}}\times U(1)_{\text{V}}\times U(1)_{\text{C}}$ as we have discussed in this work.
Symmetry of $\mathcal{H}_{0}$
First of all, let us briefly review the BM and THF models. In these models we have $\{H_{0},C_{2z}\tilde{P}\}=0$, and in the chiral limit, $\{H_{0},\tilde{\sigma}_{z}\tau_{z}\}=[H_{0},\tilde{R}]=0$ ($H_{0}$ is the bare dispersion of the BM or THF model; in the THF model, a modified sublattice operator similar to our $\sigma_{z}$ is used).
We see that only in the chiral limit the $\tilde{R}$ symmetry remains. As a matter of fact, the relations $\{H_{0},C_{2z}\tilde{P}\}=0$ and (approximately) $\{H_{0},\tilde{\sigma}_{z}\tau_{z}\}=0$ imply that the symmetry is broken by energy scales as large as the full bandwidth of the theory. Projecting to the flat bands effectively reduces the bandwidth and in strong coupling, $H_{\text{int}}\gg H_{0}\approx 0$, we have $[C_{2z}\tilde{P},H_{0}]=[C_{2z}\tilde{P},H_{\text{int}}]=0$ and additionally $[\tilde{\sigma}_{z}\tau_{z},H_{0}]=[\tilde{\sigma}_{z}\tau_{z},H_{\text{int}}]=0$ in the chiral limit. Nonetheless, a proper many-body projection to the flat bands will presumably induce new components in $H_{0}$ as we have shown in this work, disallowing the approximation $H_{0}\approx 0$.
Back to our system, in the normal state intervalley scattering is negligible and the spins are identical, so the $U(2)_{K}\times U(2)_{K^{\prime}}$ group is exactly preserved by the dispersion. For the $C_{2z}P$, $R$ and $\sigma_{z}\tau_{z}$ symmetries, let us focus on the $n_{B}=4$ theory. As stated in the main text, we can write ${h}(\bm{k})$ in the vortex Chern basis and decompose it into real components,
${h}(\bm{k})$ $={h}_{00}(\bm{k})+{h}_{z0}(\bm{k})\tau_{z}+{h}_{x0}(\bm{k})\sigma_{x}+{h}_{yz}(\bm{k})\sigma_{y}\tau_{z}+{h}_{y0}(\bm{k})\sigma_{y}+{h}_{xz}(\bm{k})\sigma_{x}\tau_{z},$ (S26)
The strength of the different components is summarized in Table (S2). The dominant term is $h_{x0}(\bm{k})\sigma_{x}+h_{yz}(\bm{k})\sigma_{y}\tau_{z}$, with ${h}_{00}(\bm{k})$ also sizable at the $\Gamma_{M}$ point. In fact, $h_{x0}(\bm{k})$ and $h_{yz}(\bm{k})$ are the main sources of symmetry breaking in $n_{B}=4$ MATBG, as we discuss in the main text. We have then the relations (neglecting $h_{00}$)
$\big{\{}C_{2z}P,\mathcal{H}_{0}\big{\}}\approx 0,\quad\quad\quad\big{\{}\sigma_{z}\tau_{z},\mathcal{H}_{0}\big{\}}\approx 0,\quad\quad\quad\big{[}R,\mathcal{H}_{0}\big{]}\approx 0.$ (S27)
Notice that the property of the chiral limit, $\{\sigma_{z}\tau_{z},\mathcal{H}_{0}\}\approx 0$, is realized to a good approximation in $\mathcal{H}_{0}$. The breaking of the chiral symmetry in our model occurs in ${\mathcal{H}}_{\text{int}}$ owing to $\overline{\sigma}_{z}\neq\sigma_{z}$ as we discussed above.
Then, $\mathcal{H}_{0}$ breaks the $C_{2z}P$ and $\sigma_{z}\tau_{z}$ symmetry. All the more, in theories with larger bandwidth, say $n_{B}=20$, and assuming that the relations of Eq. (S27) are approximately satisfied, the $C_{2z}P$ and $\sigma_{z}\tau_{z}$ symmetries are more seriously broken. In this respect, the authors of Ref. [68] discovered a chiral operator having a simple form in the THF basis, denoted $\tilde{S}$, with $[\tilde{S},H_{\text{int}}]\approx 0$ and $\{\tilde{S},H_{0}\}\approx 0$; the non-anticommuting terms being of the order of the flat bandwidth. This ’third chiral’ operator can be combined with the particle-hole operator resulting in $[C_{2z}\tilde{P}\tilde{S},H_{0}]\approx 0$ and $[C_{2z}\tilde{P}\tilde{S},H_{\text{int}}]\approx 0$. It was shown that $\tilde{S}$ is consistent with the identity in (most of) the flat bands, resulting in that $C_{2z}\tilde{P}\tilde{S}$ approximately anticommutes with the flat bands and approximately commutes with the higher bands of $H_{0}$. This is in contrast to $C_{2z}\tilde{P}$ that anticommutes with all the bands. Similarly, the operator $\tilde{\sigma}_{z}\tau_{z}\tilde{S}$ approximately commutes with $H_{\text{int}}$ and with the high-energy bands of $H_{0}$ (even though the relation $\{\tilde{\sigma}_{z},H_{0}\}\approx 0$ is expected to become worse on the higher bands), and approximately anticommutes with the flat bands.
We anticipate that the third chiral operator $\tilde{S}$ of Ref. [68] will lead to analogous relations for the operators $C_{2z}PS$, $\sigma_{z}\tau_{z}S$, ${\mathcal{H}}_{\text{int}}$, $\mathcal{H}_{0}$ of the projected theory, effectively replacing $\sigma_{z}\tau_{z}$ and $C_{2z}P$ by $C_{2z}PS$ and $\sigma_{z}\tau_{z}S$ as the generators of the approximate symmetries.
bare ($\epsilon_{r}^{-1}=0$, $U=0$ eV) $\epsilon_{r}=50$, $U=0.5$ eV $\epsilon_{r}=10$, $U=4$ eV
minimal $\overline{\sigma}_{z}$ eigenvalue $0.524$ $0.548$ $0.638$
maximal $\overline{\sigma}_{z}$ eigenvalue $0.849$ $0.858$ $0.890$
 minimal $C_{2z}\overline{P}$ singular value $0.975$ $0.978$ $0.988$
maximal $C_{2z}\overline{P}$ singular value $0.996$ $0.997$ $0.998$
avBZ $|{h}_{0z}|$ (meV) $0.16$ $0.27$ $0.76$
avBZ $|{h}_{y0}+i{h}_{xz}|$ (meV) $0.35$ $0.39$ $0.43$
avBZ $|{h}_{x0}+i{h}_{yz}|$ (meV) $0.64$ $1.36$ $5.85$
$|{h}_{x0}(\Gamma_{M})+i{h}_{yz}(\Gamma_{M})|$ (meV) $4.34$ $7.04$ $20.99$
avBZ ${h}_{00}$ (meV) $0.54$ $0.69$ $1.23$
${h}_{00}(\Gamma_{M})$ (meV) $5.79$ $6.45$ $8.21$
Table S2: Relevant figures for the breaking of the $U(4)\times U(4)$ symmetry in the $n_{B}=4$ system. Singular values of $C_{2z}\overline{P}$ different from $1$ break the flat symmetry of $\mathcal{H}_{\text{int}}$, and eigenvalues of $\overline{\sigma}_{z}$ different from $1$ break the chiral symmetry of $\mathcal{H}_{\text{int}}$. The components $h_{0z}(\bm{k})$, $h_{x0}(\bm{k})$ and $h_{yz}(\bm{k})$ break the flat symmetry of $\mathcal{H}_{0}$, and the components $h_{0z}(\bm{k})$, $h_{y0}(\bm{k})$ and $h_{xz}(\bm{k})$ break the chiral symmetry of $\mathcal{H}_{0}$. avBZ denotes the average over the Brillouin zone.
V Topological phase transition in the graphene subtraction scheme
It has been argued that the tight-binding function $t(\bm{r})$ is derived in such a way that some long range exchange is already taken into account, so one must subtract a counter-term to the interaction in order to avoid double counting. In the graphene subtraction scheme, the Hartree-Fock potential of two decoupled graphene layers at neutrality is chosen for subtraction. In other words, $H_{\text{int}}$ is normal-ordered [90] with respect to the state of two decoupled graphene layers at the neutrality point. The Hamiltonian then reads
${H}=H_{\text{TB}}+{H}_{\text{int}}$ (S28)
$H_{\text{TB}}=\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}$ (S29)
${H}_{\text{int}}=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}}ss^{\prime}}V(\bm{r}-\bm{r^{\prime}}):c^{\dagger}_{\bm{r}s}c_{\bm{r}s}c^{\dagger}_{\bm{r^{\prime}}s^{\prime}}c_{\bm{r^{\prime}}s^{\prime}}:+\sum_{\bm{r}}U:c^{\dagger}_{\bm{r}\uparrow}c_{\bm{r}\uparrow}c^{\dagger}_{\bm{r}\downarrow}c_{\bm{r}\downarrow}:$
$\hskip 19.0633pt=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}}ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r}s}c^{\dagger}_{\bm{r^{\prime}}s^{\prime}}c_{\bm{r^{\prime}}s^{\prime}}+\sum_{\bm{r}}Uc^{\dagger}_{\bm{r}\uparrow}c_{\bm{r}\uparrow}c^{\dagger}_{\bm{r}\downarrow}c_{\bm{r}\downarrow}$
$\hskip 29.87547pt-\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\langle c^{\dagger}_{\bm{r}s^{\prime}}c_{\bm{r}s^{\prime}}\rangle_{\text{gr}}\ c^{\dagger}_{\bm{r}s}c_{\bm{r}s}+\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r})\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{\text{gr}}\ c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}$
$\hskip 29.87547pt-\sum_{\bm{r}s}U\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{\text{gr}}\ c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant},$ (S30)
with $\langle...\rangle_{\text{gr}}$ the expectation value in the state of the decoupled layers of graphene. The mean-field Hamiltonian reads
${H}_{\text{MF}}=$ $\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\bigg{(}\langle c^{\dagger}_{\bm{r}s^{\prime}}c_{\bm{r}s^{\prime}}\rangle_{0}-\langle c^{\dagger}_{\bm{r}s^{\prime}}c_{\bm{r}s^{\prime}}\rangle_{\text{gr}}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}$
$-\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r})\bigg{(}\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{0}-\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{\text{gr}}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\sum_{\bm{r}s}U\bigg{(}\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{0}-\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{\text{gr}}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant}.$ (S31)
It was assumed that the contributions to the Fock matrix $\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{0}$ coming from high-energy states are essentially identical to those of the decoupled layers, contributing to $\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{\text{gr}}$. Hence, the virtual processes with remote bands would be suppressed and the normal state will be close to the ground state of the tight-binding Hamiltonian.
Contrary to this expectation, our calculations with the graphene subtraction also display significant band widening, as depicted in Fig. S3. All the more, for some value of $\epsilon_{r}$ between $50$ and $10$ we find a topological phase transition moving the one-dimensional (two-dimensional) representations (per valley) at $\Gamma_{M}$ from the flat (remote) bands to the remote (flat) bands; the spectral weight redistribution in the graphene subtraction scheme is severe. Notice that the flat bands become entangled with the higher bands making the $n_{B}=4$ projector ill-defined.
Figure S3: Band structures of the symmetric state in the graphene subtraction scheme for $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0,5$ eV, compared to the non-interacting (bare) bands. The system underwent a topological phase transition between $\epsilon_{r}=50$ and $10$ and the flat bands become doubly (per valley) degenerate at $\Gamma_{M}$.
VI Additional results: many-body projection
We have described how our many-body projection algorithm takes into account interacting effects from the integrated-out states. Then a non-projection, this is, a projection onto all bands, must return the original tight-binding system, $\mathcal{H}=H$, with the only difference that in $\mathcal{H}$ the density is measured with respect to $\rho_{0\bm{r}}$ (Eqs. (S12),(S13)) and in $H$ with respect to the uniform $1/2$ background (Eq. (S2)). $\mathcal{H}_{0}$ will then absorb the discrepancies in the definitions of the interaction. However, these differences are essentially negligible, as we can see in Fig. S4. This is consistent with the approximate relation $\rho_{0\bm{r}}=\sum_{\bm{k},n<0}|\langle\bm{r}|\bm{k}n\rangle|^{2}\approx\frac{1}{2}\sum_{\bm{k}n}|\langle\bm{r}|\bm{k}n\rangle|^{2}=\frac{1}{2}$.
Figure S4: Band structure of $\mathcal{H}_{0}$ in the non-projection ($n_{B}=11908$) model for $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0,5$ eV, compared to the non-interacting (bare) bands. They are almost identical.
Figure S5: (a) Bare tight-binding, symmetric state and $\mathcal{H}_{0}$ ($n_{B}=4,20$) bands for coupling strength $\epsilon_{r}=50$, $U=0.5$ eV. Due to the weak coupling, the renormalized bands are only slightly wider. (b-c) The components of $h(\bm{k})$ of the flat-band theory ($n_{B}=4$) with (b) $\epsilon_{r}^{-1}=0$, $U=0$ eV (bare) and (c) $\epsilon_{r}=50$, $U=0.5$ eV.
VII Additional results: correlated states
Assuming preserved moiré translations, the self-consistent states can be described by the one-particle density matrix, $P(\bm{k})_{\alpha\beta}=\frac{1}{2}\Big{(}\mathbb{1}_{\alpha\beta}+Q(\bm{k})_{\alpha\beta}\Big{)}=\langle c^{\dagger}_{\bm{k}\alpha}c_{\bm{k}\beta}\rangle,$ with the properties $P(\bm{k})^{2}=P(\bm{k})=P(\bm{k})^{\dagger}$ and $\text{tr}(P(\bm{k}))=$ the number of occupied electrons at $\bm{k}$. The indices $\alpha,\beta=(\tau\sigma s)$ denote the valley-sublattice-spin flavor and delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the expectation value. For spin-singlets and 4 occupied electrons at each $\bm{k}$, $Q(\bm{k})$ can be decomposed into products of $\sigma$ and $\tau$ matrices with real coefficients, $Q(\bm{k})=\sum_{ij}\langle\sigma_{i}\tau_{j}(\bm{k})\rangle\sigma_{i}\tau_{j}$. Under a unitary transformation $\mathcal{U}_{n}(\varphi)=\text{exp}(i\varphi\mathcal{G}_{n}/2)$, with $\mathcal{G}=\sum_{\bm{k}}c^{\dagger}_{\bm{k}}Gc_{\bm{k}}$, the $Q(\bm{k})$ matrix transforms to $U^{*}_{n}(\varphi)Q(\bm{k})U^{T}_{n}(\varphi)$ with $U_{n}(\varphi)=\text{exp}(i\varphi G_{n}/2)$. We obtained self-consistent Hartree-Fock states in the flat-band ($n_{B}=4$) system on a $24\times 24$ momentum grid; the band structures were computed afterwards on a finer grid.
In Fig. S6 we study the KIVC (initial seed $Q(\bm{k})=\sigma_{y}\tau_{y}$ on the self-consistency loop), OP ($Q(\bm{k})=\sigma_{z}$), VP ($Q(\bm{k})=\tau_{z}$) and TIVC ($Q(\bm{k})=\sigma_{x}\tau_{x}$) states. Notice that, for the initial seeds, $|\text{KIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{OP}\rangle$, $|\text{VP}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{TIVC}\rangle$, $|\text{KIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{flat}})|\text{VP}\rangle$ and $|\text{TIVC}\rangle=\text{exp}(i\pi/4\ \mathcal{G}_{\text{flat}})|\text{OP}\rangle$.
In Fig. S6(a), KIVC and OP are insulating while VP and TIVC are gapless. This can be understood because KIVC and OP break $C_{2z}\mathcal{T}$ and open a gap at the $K$ points; while VP and TIVC preserve $C_{2z}\mathcal{T}$ and the Dirac cones do not gap out but split in energy. Given that the bandwidth of $\mathcal{H}_{0}$ is comparable to the interaction scale (the splitting at $K_{M}$), the system remains metallic.
This can also be understood in a more algebraic way. The properties $[\sigma_{x}\tau_{x},\sigma_{x}]=[\tau_{z},\sigma_{x}]=[\sigma_{x}\tau_{x},\sigma_{y}\tau_{z}]=[\tau_{z},\sigma_{y}\tau_{z}]=0$ and $P(\bm{k})^{2}=P(\bm{k})$ disallow a smooth transition between the different orders ($\tau_{z}$, $\sigma_{x}\tau_{x}$ at the BZ boundaries and $\tau_{x}$, $\sigma_{y}\tau_{z}$ at the BZ center); It has to be discontinuous, i.e. through a band crossing. In the TIVC phase we find multiple additional order parameters like $\sigma_{y}$, $\sigma_{x}\tau_{z}$ and $\sigma_{y}\tau_{x}$ in a small region around the band crossing, and the transition is sharp, but smooth. The consequence is the band crossings become avoided crossing with small gaps. Moreover, the different $C_{2x}$ eigenvalues of the valence bands at $\Gamma_{M}$ and $M_{M}$ oblige a genuine crossing along the $\Gamma_{M}M_{M}$ line in the form of a Dirac cone [98]. On the other hand, the OP and KIVC states have $Q=\sigma_{z}$ and $Q=\sigma_{y}\tau_{y}$ respectively at the Brillouin zone boundaries. These matrices anticommute with $\sigma_{x}$ and $\sigma_{y}\tau_{z}$ and thus a continuous transition between the different order parameters is allowed; the self-consistent OP and KIVC states are insulators with a smooth $Q$ matrix (modulo the vortex at $\Gamma_{M}$ coming from the gauge choice).
In Fig. S5(b) we plot the main order parameters of KIVC and OP. The chiral symmetry relating the two phases is evident, as well as in Fig. S5(a) in the band structures of OP and KIVC, and VP and TIVC.
In Fig. S6(c) we consider the self consistent KIVC state under flat and chiral transformations, $\mathcal{U}_{\text{flat}}(\varphi)|\text{KIVC}\rangle=\text{exp}(i\varphi/2\ \mathcal{G}_{\text{flat}})|\text{KIVC}\rangle$ and $\mathcal{U}_{\text{chiral}}(\varphi)|\text{KIVC}\rangle=\text{exp}(i\varphi/2\ \mathcal{G}_{\text{chiral}})|\text{KIVC}\rangle$. We plot the expectation value of the interaction ${\mathcal{H}}_{\text{int}}$ and the total energy ${\mathcal{H}}$ for varying $\varphi$. This provides an alternative confirmation that $\mathcal{H}_{0}$ $(={\mathcal{H}}-{\mathcal{H}}_{\text{int}})$ induces a strong $\mathcal{G}_{\text{flat}}$ breaking, while ${\mathcal{H}}_{\text{int}}$ is approximately $U(4)\times U(4)$ invariant. On the other hand, the smaller amplitude of the ${\mathcal{H}}_{\text{int}}$ curves of the flat rotations was expected form the larger singular values of $C_{2z}\overline{P}$ (Fig S2(b) and Table S2).
In Fig. S7 we show the band structure and order parameters of the NSM, QAH and SP phases. The Dirac cones in NSM are visible in all panels, and the vortex at $\Gamma_{M}$ coming from the gauge choice is visible in the phase plot. The SP phase is metallic for the same reason as the VP.
Figure S6: (a) Band structures of the Kramers intervalley coherent (KIVC), orbital polarized (OP), time-reversal invariant intervalley coherent (TIVC) and valley polarized (VP) states, for coupling strengths $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0.5$ eV. (b) Order parameters of KIVC and OP. (c) Expectation values of the energies of the rotated states $\mathcal{U}_{\text{flat}}(\varphi)|\text{KIVC}\rangle$ and $\mathcal{U}_{\text{chiral}}(\varphi)|\text{KIVC}\rangle$. The approximate chiral symmetry is manifest in (a-c) and the broken flat symmetry in (a).
Figure S7: (a) Band structure of the valley $K$ valence band and order parameter of the nematic semimetal (NSM). (b) Band structure and order parameters of the Quantum Anomalous Hall (QAH) state. (c) Band structure of the spin polarized state (SP).